40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 9589     Entry time: Fri Jan 31 18:41:25 2014
Author: manasa 
Type: Update 
Category: General 
Subject: IFO alignment update 

[EricQ, Gabriele, Manasa]

We found we had lost the Y arm pointing from yesterday. We tried to recover the pointing for a couple of hours and finally decided to take the ETMY heavy door off.

The input beam was aligned to the Y arm. We also got AS and REFL out of vacuum and on the cameras.

We put back the light doors and tried to lock the arms, but did not succeed as yet.

Things to do:
1. Lock arms for IR
2. Realign POP path
3. Recenter all oplevs
4. Try to check the state of PRC after the length change
5. Take in-vacuum pictures

ELOG V3.1.3-